Suppose that f(2) = -3, 9(2) = 4, f(2) = -5, and g(2) = 1. Find h(2). (a) h(x) = 3f(x) - 2g(x) h(2) h(x) = f(x)g(x) (b) h
Answer #1

- = (2 یا 4 - 4 = (2 و 3- (على ا = (2 او و hanj= 3 ten- 4 رد (+3 = (h c (2 او 2 - 2 الم3x = 2 (در2 (راوه - 3 -5 a | -17 = 2 -ا (3-) - 5 - 4 = (2 * (4) 2 سه - 20 +3 )6 ا۔ كل 8 h) - 2») رط + ( ا لد - ( او ( رمل+1)= ( 5 2 (الے +1) ) = (1-3) 1 وبه 8 4 ((Thanks